# ModSec-AdvLearn: Countering Adversarial SQL Injections with Robust Machine Learning

### v3

@@ IV-C Evaluation of ModSec-Learn
Robustness against Adversarial SQLi. As shown in Fig. 2, the ModSec-Learn models suffer the presence of adversarial attacks, particularly with the WAF-A-MoLE dataset, but they still outperform the vanilla ModSecurity. Analyzing the results from the WAF-A-MoLE dataset, it is evident that, the best performance is achieved with PL4, except for the RF and SecSVM models. This may be since, in this case, the adversarial SQLi attacks are able to exploit the rules enabled by PL3 and PL4 to evade the model, by removing patterns that were considered important at training time. On the other hand, with the ModSec-Learn dataset, it is observed that starting from PL2, the detection capabilities exhibit a fairly consistent trend. While there is a slight deterioration, performance remains relatively stable across the different PLs.
@@
Figure 3: Weight values learned at PL4 by ModSec-Learn LR-$\ell_{1}$ (blue) and ModSec-Learn LR-$\ell_{2}$ (light red), and the weights used by ModSecurity (green). Rules are expressed as their IDs (all starting with 942).
@@ IV-D Evaluation of ModSec-AdvLearn
Evaluation on Clean Samples. In the absence of attack, ModSec-AdvLearn has comparable performance to ModSec-Learn, except for SecSVM which shows an improvement in TPR on the ModSec-Learn dataset (cf. the violet solid lines in Fig. 4, bottom row). The reason is that SecSVM is retrained on less pessimistic attacks when considering problem-space AT, thereby yielding an improved robustness-accuracy tradeoff.
@@
Figure 4: ROC curves of ModSec-Learn and ModSec-AdvLearn, evaluated on test/test-adv (solid/dashed lines) at PL4, for WAF-A-MoLE (top) and ModSec-Learn (bottom) datasets.
Figure 5: Top: Activation probability of CRS rules on malicious (cyan) and adversarial (orange) SQLi samples optimized against ModSec-Learn SVM with $\ell_{1}$ regularization. Rules are expressed as their IDs (all starting with 942). Bottom: Weights learned by ModSec-Learn SVM-$\ell_{1}$ (cyan) and ModSec-AdvLearn SVM-$\ell_{1}$ (orange) for the SQLi rules expressed as their IDs (all starting with 942) on WAF-A-MoLE dataset.
@@ V Related work
ModSecurity and CRS. Previous work has considered the impact of different types of web security threats on ModSecurity when using the CRS configurations [24, 25]. However, unlike ours, a very limited number of attack samples is normally used (e.g., [24] uses only 27 samples), without providing any detailed investigation of the trade-off between TPR and FPR.
@@ VI Conclusions and Future Work
In this work we proposed ModSec-AdvLearn, a novel methodology for training ML classifiers using the CRS rules as input features. This allows learning how to optimally tune the severity levels (i.e., the weights) of the CRS rules, yielding an improved trade-off between detection and false positive rates. Furthermore, our approach relies upon a novel problem-space adversarial training procedure that incorporates knowledge of state-of-the-art SQLi manipulations to counter the presence of adversarial SQLi attacks. Among the main findings, we show that ModSec-AdvLearn improves the detection rate of the vanilla ModSecurity by 30%, while removing 50% of the CRS rules through embedded feature selection with $\ell_{1}$ regularization, and improves the adversarial robustness up to 85%. Hence, we can state that our methodology provides a first, concrete example of how adversarial machine learning can be used in the problem space to effectively enhance the robustness of WAFs against adversarial attacks.
We foresee several promising avenues for advancing our work. First, although in this work we only target SQLi attacks, our methodology is general enough to tackle other web threats like cross-site scripting (XSS). Secondly, we also see future developments in evaluating other state-of-the-art ML-based WAFs. Indeed, we think that the same results can be also obtained on more advanced models such as Convolutional Neural Networks (CNN) [30], as well as on different feature representation approaches [31]. This is also true for commercial WAFs. To this end, an interesting future extension of this work is to evaluate them in terms of transferability [32] of adversarial SQLi attacks optimized on ModSecurity.
@@ Acknowledgments
This research has been partly supported by the TESTABLE project, funded by the EU H2020 research and innovation program (grant no. 101019206); the ELSA project, funded by the Horizon Europe research and innovation program (grant no. 101070617); projects FAIR (PE00000013) and SERICS (PE00000014) under the NRRP MUR program funded by the EU – NGEU. This work has been carried out while C. Scano was enrolled in the Italian National Doctorate on AI run by the Sapienza University of Rome in collaboration with the University of Cagliari.

### v4

@@ IV-C Evaluation of ModSec-Learn
Robustness against Adversarial SQLi. As shown in Fig. 2, the ModSec-Learn models suffer the presence of adversarial attacks, particularly with the WAF-A-MoLE dataset, but they still outperform the vanilla ModSecurity. Analyzing the results from the WAF-A-MoLE dataset, it is evident that, the best performance is achieved with PL4, except for the RF and SecSVM models. This may be since, in this case, the adversarial SQLi attacks are able to exploit the rules enabled by PL3 and PL4 to evade the model, by removing patterns that were considered important at training time. On the other hand, with the ModSec-Learn dataset, it is observed that starting from PL2, the detection capabilities exhibit a fairly consistent trend. While there is a slight deterioration, performance remains relatively stable across the different PLs. Among all evaluated models, SecSVM is the most robust, offering strong generalization and adversarial robustness. Even if RF was included to explore non linear alternatives, its performance is comparable to linear models like SecSVM, while worsening robustness. This confirms that linear models are sufficient to achieve excellent accuracy and are preferable given their robustness and transparency. In addition, linear models can be readily applied to the existing CRS system by updating the rule weights after model training. [^10] https://owasp.org/www-project-waf-advanced-ruleset-management/ This eases practical deployment while preserving interpretability of decisions — an important desideratum for web security.
@@
Figure 3: Weight values learned at PL4 by ModSec-Learn LR-$\ell_{1}$ (blue) and ModSec-Learn LR-$\ell_{2}$ (light red), and the weights used by ModSecurity (green). Rules are expressed as the last three digits of their IDs (all starting with 942).
@@ IV-D Evaluation of ModSec-AdvLearn
Evaluation on Clean Samples. In the absence of attack, ModSec-AdvLearn has comparable performance to ModSec-Learn, except for SecSVM which shows an improvement in TPR on the ModSec-Learn dataset (cf. the violet solid lines in Fig. 4, third row). The reason is that SecSVM is retrained on less pessimistic attacks when considering problem-space AT, thereby yielding an improved robustness-accuracy tradeoff.
@@
Figure 4: ROC curves of ModSec-Learn/AdvLearn (first/second column) on test/test-adv (solid/dashed lines), for WAF-A-MoLE/ModSec-Learn (top/bottom) datasets.
Figure 5: Top: Activation probability of CRS rules (expressed using the last three digits of their IDs, which always starts with 942) on malicious/adversarial (cyan/orange) SQLi samples optimized against ModSec-Learn SVM-$\ell_{1}$ on the WAF-A-MoLE dataset. Bottom: Rule weights learned by ModSec-Learn/ModSec-AdvLearn SVM-$\ell_{1}$ (cyan/orange) on the WAF-A-MoLE dataset.
@@ V Related work
ModSecurity and CRS. Previous work has considered the impact of different types of web security threats on ModSecurity when using the CRS configurations [24, 25]. However, unlike ours, a very limited number of attack samples is normally used (e.g., [24] uses only 27 samples), without providing any detailed investigation of the trade-off between TPR and FPR. Other approaches [26, 27] have applied ML to detect web threats and only used ModSecurity as a baseline for comparison, without even evaluating adversarial robustness.
@@ VI Conclusions and Future Work
In this work, we proposed ModSec-AdvLearn, a novel methodology for training ML classifiers using the CRS rules as input features. This allows learning how to optimally tune the severity levels (i.e., the weights) of the CRS rules, yielding an improved trade-off between detection and false positive rates. Furthermore, our approach relies upon a novel problem-space adversarial training procedure that incorporates knowledge of state-of-the-art SQLi manipulations to counter the presence of adversarial SQLi attacks. Among the main findings, we show that ModSec-AdvLearn improves the detection rate of the vanilla ModSecurity by 30%, while removing 50% of the CRS rules through embedded feature selection with $\ell_{1}$ regularization. It also improves adversarial robustness up to 85% via robust linear models, without hindering interpretability of decisions and providing ease of integration with the current CRS implementations. We can thus state that our methodology provides a first, concrete example of how adversarial machine learning can be used to effectively enhance the robustness of WAFs against adversarial attacks, highlighting novel, promising directions towards designing robust machine learning models for cybersecurity-related applications.
We foresee several other promising avenues for advancing our work. First, although in this work we only target SQLi attacks, our methodology is general enough to tackle other web threats like cross-site scripting (XSS). In this direction, future work could also explore the integration of automated pentesting tools that leverage large language models [32], to extend the capabilities of WAF-A-MoLE. Second, we also see future developments in evaluating other state-of-the-art ML-based WAFs. Indeed, we think that the same results can also be obtained on more advanced models such as Convolutional Neural Networks (CNN) [33], as well as on different feature representation approaches [34, 27]. This is also true for commercial WAFs. To this end, an interesting future extension of this work is to evaluate them in terms of transferability [35] of adversarial SQLi attacks optimized on ModSecurity.
@@ Acknowledgments
This research has been partly supported by the TESTABLE project, funded by the EU H2020 research and innovation program (grant no. 101019206); the ELSA project, funded by the Horizon Europe research and innovation program (grant no. 101070617); projects FAIR (PE00000013) and SERICS (PE00000014) under the NRRP MUR program funded by the EU – NGEU. This work was carried out while C. Scano was enrolled in the Italian National Doctorate on AI run by the Sapienza University of Rome in collaboration with the University of Cagliari.